OpenStudy (anonymous):

After reading to a kidergarten class in library, Tobey gives the stickers out at random, Hes has 9 zoo animals stickers and 16 strach ans sinff stickers.what is the probability that the first child will get a starch and sinfo ticker and the second child get a zoo animal sticker?

OpenStudy (anonymous):

Please help

OpenStudy (studygurl14):

Hello @Janelle01 ! Welcome to Openstudy!

OpenStudy (anonymous):

Thx

OpenStudy (igreen):

First, find the total amount of stickers.

OpenStudy (igreen):

9 + 16 = ?

OpenStudy (anonymous):

25

OpenStudy (studygurl14):

Correct. @Janelle01 To find the probability of someone getting a scratch-n-sniff, you take the amount of scratch-n-sniff stickers and divide it by the total amount of stickers.

OpenStudy (anonymous):

16/25=64%

OpenStudy (studygurl14):

Correct. Nice job.

OpenStudy (studygurl14):

Now, the next part.

OpenStudy (anonymous):

So what about the zoo animals stickers

OpenStudy (studygurl14):

Take the amount of zoo animal stickers and divide it by the total amount of stickers.

OpenStudy (anonymous):

9/25=36%

OpenStudy (studygurl14):

Correct. Now we're finding the probability of both events occurring correct? That means we have to multiply those two probabilities together.

OpenStudy (anonymous):

Ok so .64×.34 right

OpenStudy (studygurl14):

yep.

OpenStudy (anonymous):

.2304

OpenStudy (studygurl14):

Yes. Depending on what they want as your answer, you might have to convert to percentage form

OpenStudy (anonymous):

So 23 and 1/25

OpenStudy (studygurl14):

No, to get a percent, you multiply by 100 and add the % sign

OpenStudy (anonymous):

So .2304 × 100= 23 1/25%

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

U their